### v1

@@ 3 Discussion
We present a framework for assessing LLM performance for real-world medical tasks. Our clinician-validated taxonomy provides a structure for summarizing models’ strengths and limitations across medical tasks. High clinician agreement (96.7%) in assigning subcategories to categories suggests the taxonomy effectively captures how healthcare professionals conceptualize their work. Our benchmark suite reveals nuances in model capabilities that are not seen with medical knowledge benchmarks alone. The higher performance in communication tasks than administrative ones may stem from administrative workflows using data not seen during training, warranting caution in healthcare AI implementation for back-office tasks without quantifying task-specific performance. This framework addresses the primary limitations of current benchmarks, such as the lack of using real-world data, use of evaluation setups that do not match real-world settings, and limited task diversity. The cost-performance trade-off shows that while reasoning models have superior performance, their substantially higher costs may not justify their deployment for all tasks. In a resource-constrained setting, models such as Claude 3.5 Sonnet offer a balance, achieving a win-rate of 0.63 at a lower cost. Finally, our LLM-jury based approach addresses a critical gap in current evaluation approaches. By beating clinician-clinician agreement, this approach enables scalable evaluation of open-ended model outputs without requiring extensive clinician time — a scarce and expensive resource.
@@
Several limitations remain. While our LLM-jury approach was validated on only two benchmarks, expanding clinician annotations across more benchmarks would strengthen the clinician–LLM agreement estimates. Our current rubrics operate at the benchmark level, but instance-level rubrics could provide better evaluation, particularly for subjective or context-dependent medical tasks where gold standard responses may not exist. Such approaches would further scale LLM-jury evaluation beyond reliance on gold standard responses [11]. Administration & Workflow emerged as the weakest performance area for all models. Understanding the underlying causes of this poor performance, whether stemming from training data limitations, task complexity, or distributional shifts, is essential for safe deployment in healthcare operations.
@@
In conclusion, MedHELM provides a comprehensive framework for assessing LLM performance across real-world medical tasks through our clinician-validated taxonomy (96.7% agreement) spanning 5 categories, 22 subcategories, and 121 tasks. Our benchmark suite of 35 datasets reveals that most models perform best in Clinical Note Generation (0.73-0.85, with one outlier at 0.56) and Patient Communication & Education (0.78-0.83), moderately in Medical Research Assistance (0.65-0.75), and worst in Clinical Decision Support (0.56-0.72) and Administration & Workflow (0.53-0.63). Reasoning models DeepSeek R1 and o3-mini led overall with win-rates of 0.66 and 0.64 respectively, though Claude models offer competitive performance (win-rate of $\sim 0.63$) at lower computational cost. Through our public leaderboard [^5] https://crfm.stanford.edu/helm/medhelm/latest/ and shared codebase [^6] https://github.com/stanford-crfm/helm, MedHELM establishes infrastructure for ongoing collaborative assessment as models evolve, advancing medical AI evaluation that better reflects the complexity of real-world medical practice.
@@ 4.1.1 Initial drafting
We based our taxonomy on tasks identified in a JAMA review [7]. Working with a clinician (MK), we reorganized these tasks into functional themes that reflect real-world activities, resulting in 98 distinct tasks organized into 21 subcategories within 5 categories:
@@ 4.3.4 Cost–Performance Analysis
We tracked token usage for both benchmark prompts and jury evaluations across all models by counting input tokens and assuming the maximum allowed output tokens. This approach was necessary because, for o3-mini, it is not possible to fully observe the token usage as the thinking tokens don’t appear in the final output. As a result, our cost estimates represent an upper bound. Using published per-1M-token pricing as of 05/12/2025, we estimated total cost per model. Cost–performance trade-offs were visualized by plotting mean win-rate against estimated total evaluation cost, highlighting models that optimize accuracy per dollar spent.
@@ Appendix C List of Benchmarks in MedHELM
Clinical Note Generation DischargeMe DischargeMe is a discharge instruction generation dataset and brief hospital course generation dataset collected from MIMIC-IV data, considering only the discharge text as well as the radiology report text. Gated Existing
@@
Patient Communication and Education MedicationQA Open text question-answer pairs regarding consumer health questions about medication. Public Existing
@@
EHRSQL Given a natural language instruction, generate an SQL query that would be used in clinical research. Public Existing

### v2

@@ 3 Discussion
We present a framework for assessing LLM performance for real-world medical tasks. Our clinician-validated taxonomy provides a structure for summarizing models’ strengths and limitations across medical tasks. High clinician agreement (96.7%) in assigning subcategories to categories suggests the taxonomy effectively captures how healthcare professionals conceptualize their work. Our benchmark suite reveals nuances in model capabilities that are not seen with current medical knowledge benchmarks alone. The higher performance in communication tasks than administrative ones may stem from administrative workflows using data not seen during training, warranting caution in healthcare AI implementation for back-office tasks without quantifying task-specific performance. This framework addresses the primary limitations of current benchmarks, such as the lack of using real-world data, use of evaluation setups that do not match real-world settings, and limited task diversity. The cost-performance trade-off shows that while reasoning models have superior performance, their substantially higher costs may not justify their deployment for all tasks. In a resource-constrained setting, models such as Claude 3.5 Sonnet offer a balance, achieving a win-rate of 0.63 at a lower cost. Finally, our LLM-jury based approach addresses a critical gap in current evaluation approaches. By beating clinician-clinician agreement, this approach enables scalable evaluation of open-ended model outputs without requiring extensive clinician time, a scarce and expensive resource.
@@
Several limitations remain. While our LLM-jury approach was validated on only two benchmarks, expanding clinician annotations across more benchmarks would strengthen the clinician–LLM agreement estimates. In addition, the uneven distribution of benchmarks across subcategories (15 of 22 contain only one benchmark) limits our ability to draw robust performance conclusions in underrepresented areas. Moreover, our current rubrics operate at the benchmark level, but instance-level rubrics could provide better evaluation, particularly for subjective or context-dependent medical tasks where gold standard responses may not exist. Such approaches would further scale LLM-jury evaluation beyond reliance on gold standard responses [12]. Administration & Workflow emerged as the weakest performance area for all models. Understanding the underlying causes of this poor performance, whether stemming from training data limitations, task complexity, or distributional shifts, is essential for safe deployment in healthcare operations.
@@
In conclusion, MedHELM provides a comprehensive framework for assessing LLM performance across real-world medical tasks through our clinician-validated taxonomy spanning five categories, 22 subcategories, and 121 tasks. Our benchmark suite of 35 datasets reveals that most models perform best in Clinical Note Generation (0.74-0.85) and Patient Communication & Education (0.76-0.89), moderately in Medical Research Assistance (0.65-0.75) and Clinical Decision Support (0.61-0.76), and worst in Administration & Workflow (0.53-0.63). Reasoning models DeepSeek R1 and o3-mini led overall with win-rates of 0.66 and 0.64, respectively, though Claude models offer competitive performance (win-rate of $\sim 0.63$) at lower computational cost. Through our public leaderboard [^5] https://crfm.stanford.edu/helm/medhelm/latest/ and shared codebase [^6] https://github.com/stanford-crfm/helm, MedHELM establishes infrastructure for ongoing collaborative assessment as models evolve, advancing medical AI evaluation that better reflects the complexity of real-world medical practice.
@@ 4.1.1 Initial drafting
We based our taxonomy on tasks identified in a JAMA review [8]. Working with a clinician (MK), we reorganized these tasks into functional themes that reflect real-world activities, resulting in 98 distinct tasks organized into 21 subcategories within five categories:
@@ 4.3.4 Cost–Performance Analysis
We tracked token usage for both benchmark prompts and jury evaluations across all models by counting input tokens and assuming the maximum allowed output tokens. This approach was necessary because, for o3-mini, it is not possible to fully observe the token usage as the thinking tokens don’t appear in the final output. As a result, our cost estimates represent an upper bound. Using published per-1M-token pricing as of 05/12/2025, we estimated the total cost per model. Cost–performance trade-offs were visualized by plotting mean win-rate against estimated total evaluation cost, highlighting models that optimize accuracy per dollar spent.
@@ Appendix C List of Benchmarks in MedHELM
Clinical Note Generation DischargeMe A dataset that provides discharge text as well as the radiology report text collected from MIMIC-IV data such that models can generate discharge instructions and brief hospital course information generation. Gated Existing
@@
Patient Communication and Education MedicationQA A dataset containing open text question-answer pairs regarding consumer health questions about medication. Public Existing
@@
EHRSQL A dataset that generates an SQL query that would be used in clinical research given a natural language instruction. Public Existing